Required information

[The following information applies to the questions displayed below.]

Shown below is information relating to the stockholders' equity of Brookdale Corporation at December 31, 2018:

11% cumulative preferred stock, $130 par, 100,000 shares authorized,
10,000 shares issued
$ 1,300,000
Common stock, $1.25 par, 1,000,000 shares authorized, 600,000 shares issued
(of which 6,000 are held in treasury)
750,000
Additional paid-in capital: preferred stock 500,000
Additional paid-in capital: common stock 900,000
Additional paid-in capital: treasury stock transactions 6,000
Treasury stock (at cost: 6,000 common shares) (192,000 )
Retained earnings 1,350,000

What was the average issue price per share of common stock?
Answer #1

Ans:

Average Issue Price of Common Stock:

1)No of Common Stock shares issued * Par Value=600,000-6000*1.25=$742,500
2)Add:Paid in Capital in excess of par value= $900,000
3) Totals (1+2)= 1,642,500
4)No of Common Stock shares issued (600,000-6000=594,000) 594,000

5) Average issue price of common Stock (3/4)-1642500/594,000= 2.765 or 2.77 (if rounded off)